# Chinese Academy of Sciences, Beijing 100190, China

### v1

@@ 2.1 Stress-tensor supermultiplet and the form factor
$-\theta_{\alpha}^{+a}\epsilon^{\alpha\beta}\theta_{\beta}^{b}{\rm tr}\Big{(}\psi^{+\gamma}_{(a}\psi^{+}_{b)\gamma}-g\sqrt{2}[\phi^{+C}_{(a},\bar{\phi}_{C\,+b)}]\phi^{++}\Big{)}$
$-{4\over 3}(\theta^{+})^{3\,a}_{\alpha}{\rm tr}\Big{(}F_{\beta}^{\alpha}\psi_{a}^{+\beta}+ig[\phi_{a}^{+B},\bar{\phi}_{BC}]\psi^{C\alpha}\Big{)}+{1\over 3}(\theta^{+})^{4}\,{\cal L}(x)\ .$ (2)
@@ 2.2 Unitarity-cut method
Hence, the supersymmetry is preserved in this scheme. [^1] We recall that in the four-dimensional cuts, the helicity sum can be performed by integrating over the Grassmann numbers of internal lines $\sum_{\textrm{physical states of}\ \ell}\xrightarrow{4-\text{dim}}\int\text{d}^{4}\eta_{\ell}$.
@@ 3.2 $D$-dimensional unitarity cuts
The $q^{2}$-cut channel shown in Figure 1(a) with external states as $(1^{+},2^{+},3^{\phi_{12}},4^{\phi_{12}})$ is given by
@@
The $q^{2}$-cut channel involves only the cut scalar states. A more complicated case is the $s_{12}$-cut channel shown in Figure 1(b), which contains gluonic and fermionic contributions and provides additional checks for our result. We consider the external states as $\{1^{+},2^{\phi_{12}},3^{\phi_{12}},4^{+}\}$, then the cut result is
@@
where the superscript $\Phi_{i}$ denotes the physical state associated with $p_{i}$. The possible physical states $(\Phi_{5},\Phi_{6})$ are $(g,\phi_{34})$, $(\phi_{34},g)$, $(\psi_{134},\psi_{234})$, and $(\psi_{234},\psi_{134})$. The most complicated terms are $(\psi_{134},\psi_{234})$, and $(\psi_{234},\psi_{134})$ that involve the fermionic tree-level results. In this case, the two tree-level building blocks can be also computed by Feynman diagrams to capture the $D$-dimensional information, as
@@
where $p_{56}$ denotes $p_{5}+p_{6}$, and the $D$-dimensional gamma traces ${\bf tr}$ defined in (18) keep the $D$-dimensional information of the loop momenta, as we discussed in Section 2. We mention here that it is practically convenient to compare them with the cut integrand numerically.222 We give a brief comment on the numerical method for $D$-dimensional kinematics. Any massless momentum can be expressed as $\ell=(\ell^{0},\ell^{0}\hat{n}_{D-1})$, where $\ell^{0}$ is the zeroth component, and $\hat{n}_{D-1}=(\cos\theta_{1}\,,\ \ldots\,,\ \sin\theta_{1}\cdots\sin\theta_{D-2})$ is a $(D-1)$-dimensional unit vector. We consider the $D>4$ to capture the information beyond the four-dimension. This form takes the advantage that the momentum conservation equations and on-shell conditions are linear for $\ell^{0}$ (except the three-point case). Then it is convenient to apply the numerical evaluation with specific dimensions such as $D=6$, which is enough for the two-loop four-point case.
@@ 4.1 Summary of the two-loop integrand
$=$ $s_{123}\left[s_{12}\text{tr}_{-}(1,\ell_{7},4,3)-s_{12\ell_{1}}\text{tr}_{-}(1,\ell_{7},4,3)+\text{tr}_{-}(1,\ell_{7},4,3,\ell_{1},2)\right]$
@@ 4.2 More details in the unitarity cuts
This unitarity cut has only contribution from $N_{1}$, since only the topology in Figure 2(a) contributes to this cut.666 The integrand based on four-dimensional cuts has been considered in Gopalka:2023doh . There seems no contribution of topology associated with $N_{1}$ in the result of Gopalka:2023doh , which would be inconsistent with the cut of Figure 3 even in the four-dimension. The $N_{1}$ part has the non-trivial contribution in the IBP reduced form and to the remainder function as will be considered in the next section.
@@
The $D$-dimensional version of the unitarity cut in Figure 3 is
@@
The $D$-dimensional unitarity cut as shown as Figure 3 involves the two-gluons cut states as
${\cal F}^{(2)}_{4}\big{|}_{\text{cut-(b)}}=$$\sum_{\{\Phi_{i}\}}{\cal A}_{6}^{(0)}(1^{+},2^{+},5^{\Phi_{5}},6^{\Phi_{6}}){\cal A}_{4}^{(0)}(3^{\phi_{12}},4^{\phi_{12}},\bar{7}^{\Phi_{\bar{7}}},\bar{8}^{\Phi_{\bar{8}}}){\cal F}_{2}^{(0)}(\bar{6}^{\Phi_{\bar{6}}},\bar{5}^{\Phi_{\bar{5}}},8^{\Phi_{8}},7^{\Phi_{7}})$
$=$ $\sum_{g}{\cal A}_{6}^{(0)}(1^{+},2^{+},5^{g},6^{g}){\cal A}_{4}^{(0)}(3^{\phi_{12}},4^{\phi_{12}},\bar{7}^{\phi_{34}},\bar{8}^{\phi_{34}}){\cal F}_{2}^{(0)}(\bar{6}^{g},\bar{5}^{g},8^{\phi_{12}},7^{\phi_{12}})\,.$ (41)
Note that the external states of the cut legs for $p_{5}$ and $p_{6}$ must be gluon states, and then the cut legs for $p_{7}$ and $p_{8}$ can only be scalar states $\phi_{12}$.
Another two-double-cut example similar to the one-loop case is shown in Figure 3
@@
Finally, the triple cut will provide the most stringent crosscheck for the integrand result. We find that all triple cuts pass directly given the result derived by the two-double cuts. Consider for example the cut shown in Figure 3:
@@ 5.1 The IBP reduction of the form factor
$\mathbb{I}_{i}=\sum_{j}c_{ij}\mathbb{I}_{ij}^{\text{IBP-input}}\,,$ (45)
where $\mathbb{I}_{ij}^{\text{IBP-input}}$ are the IBP-input integrals, defined as
$\mathbb{I}_{ij}^{\text{IBP-input}}=\prod_{k}D_{i,k}^{-\alpha_{ij,k}}\,,$ (46)
where $D_{i,k}$ is the $k$-th propagator of $i$-th topology.
@@
The IBP-input integrals are ready to be reduced to a finite set of master integrals by using IBP relations. We perform the IBP reduction by the program Kira Maierhofer:2017gsa ; Lange:2021edb . The coefficients of master integrals in the result can be simplified significantly for the half-BPS form factor if choosing the master integrals to have a uniform transcendental degree. The related UT master integrals, which will be denoted by $I^{(\ell),\text{UT}}$, are given in Abreu:2020jxa ; Canko:2020ylt ; Abreu:2021smk ; Abreu:2023rco .
@@
The parity-odd part is defined by the terms that are proportional to $\varepsilon(1234)$. Note that the $\varepsilon(1234)$ variable appears in both the factor $B$ and the definition of the parity-odd integral $I_{\text{Pen}}^{(1)}$. We mention that $I_{\text{Pen}}^{(1)}$ contributes to the $\mathcal{O}(\epsilon)$ order, and so does ${\cal G}_{2}^{(1)}$ and ${\cal I}_{{\cal F},4}^{(1)}\big{|}_{\text{odd}}$.
@@ 5.2 Two-loop remainder function
where $i,j=1,2,3,4$ and note that only five of them are independent. Not only this, the remainder function in terms of the Chen iterated path integrals contains only 34 function letters $\{m_{i}\}$ which are the same as the symbol letters. We have checked that our result is consistent with the symbol expression of the same form factor obtained via symbol bootstrap in Dixon:2022xqh .
@@ 5.3 Evaluating the remainder function
Figure 8: Evaluating the remainder function for a physical scattering process. For the figures in the first row, the lower-left linear boundary and the upper-right curve boundary correspond separately to the outgoing angle $\theta=0$ and $\theta=\pi$. The two boundaries insert symmetrically at $(0,1-m_{q}^{2})$ and $(1-m_{q}^{2},0)$ correspond separately to the soft limit with $p_{1}\to 0$ and $p_{2}\to 0$, where $m_{q}=1/10$.
@@
$s_{12}=P^{2}=10^{-16}\,,\quad s_{13}=-\frac{1}{6}\,,\quad s_{14}=-\frac{1}{18}\,,\quad s_{23}=-\frac{1}{3}\,,\quad s_{24}=-\frac{1}{9}\,,\quad s_{34}=\frac{5}{3}\,.$ (70)
Then the form factor has a finite limit at this kinematics point as
[…]
where the kinematics of ${\cal R}_{{\cal F},3}$ are
$s_{12}^{\prime}=-1/2\,,\quad s_{23}^{\prime}=5/3\,,,\quad s_{13}^{\prime}=-1/6\,,$ (72)
with $p_{1}^{\prime}=p_{1}+p_{2}$, $p_{2}^{\prime}=p_{3}$ and $p_{3}^{\prime}=p_{4}$. Moreover, in Figure 8, the remainder function at the lower-left linear boundary corresponds to $\theta=0$. We have checked that this is consistent with taking the collinear limit $p_{1}\parallel p_{2}$, and the limit result is
${\cal R}_{{\cal F},4}^{(2)}\xrightarrow{\theta\to 0}{\cal R}_{{\cal F},3}^{(2)}\approx-5.398066341803370-3.949636854179309i\,,$ (73)
where the kinematics of ${\cal R}_{{\cal F},3}^{(2)}$ are
@@
with $p_{1}^{\prime}=p_{1}+p_{2}$, $p_{2}^{\prime}=p_{3}$ and $p_{3}^{\prime}=p_{4}$.
@@
The lightlike form factor (with $q^{2}=0$) for the stress-tensor supermultiplet has been obtained using the master-integral bootstrap method in Guo:2022qgv , which is shown to have the directional dual conformal symmetry and depends on only the three conformal ratios
@@
where $\delta\to 0$ and which satisfies $q^{2}=\delta$. Then the form factor has a finite limit at this kinematics point as
[…]
where the superscript ${\rm LL}$ denotes lightlike. In the next section, we will show that the symmetry emerges also at the integrand level.
@@ 5.4 Triple-collinear limit
$u_{2}=\left[1+T^{2}+\frac{S^{2}[S_{2}T_{2}(1+f_{2}^{2})+f_{2}(1+S_{2}^{2}+T^{2}+T_{2}^{2})]}{f_{2}S_{2}^{2}}\right]^{-1}\,,$ (79)
@@
To take the triple-collinear limit in the region related to the process $34\to 12(-q)$, the corresponding amplitude result should be obtained in the physical region related to the process ${\hat{1}}{\hat{4}}\to{\hat{2}}{\hat{3}}{\hat{5}}{\hat{6}}$. The correspondence may be understood from a symmetry point of view: the amplitude has only the flipping symmetry $\hat{1}\leftrightarrow\hat{4}$ in the physical region just as the form factor has only the flip symmetry $1\leftrightarrow 3$. The correspondence of the triple-collinear limits of the form factor and the amplitude can be illustrated in Figure 9. The triple-collinear limit $p_{1}\parallel p_{2}\parallel p_{3}$ of the four-point form factor is related to the triple-collinear limit $\hat{p}_{1}\parallel\hat{p}_{2}\parallel\hat{p}_{3}$ of the six-gluon amplitude. [^8] The form factor has two equivalent triple-collinear limits by taking $p_{1}\parallel p_{2}\parallel p_{3}$ and $p_{4}\parallel p_{1}\parallel p_{2}$, where $p_{1}$ and $p_{2}$ are out-going momenta. The amplitude has two similar limits by taking $\hat{p}_{1}\parallel\hat{p}_{2}\parallel\hat{p}_{3}$ and $\hat{p}_{2}\parallel\hat{p}_{3}\parallel\hat{p}_{4}$, where $\hat{p}_{2}$ and $\hat{p}_{3}$ are out-going momenta.
Below we check the relation (84).
@@
It is notable that the full analytic remainder ${\cal R}_{{\cal A},6}^{(2)}$ in the physical region ${\hat{1}}{\hat{4}}\to{\hat{2}}{\hat{3}}{\hat{5}}{\hat{6}}$ is still not yet available in the literature. To apply our check, we note that the $2\rightarrow 4$ amplitude has been studied in a singular configuration, which corresponds to the limit in which a hexagonal Wilson loop develops a self-crossing Dixon:2016epj . In this self-crossing limit, the cross ratios are highly constrained, for example, for the process ${\hat{1}}{\hat{4}}\to{\hat{2}}{\hat{3}}{\hat{5}}{\hat{6}}$ one has
@@
Then the Mandelstam variables for ${\cal R}_{{\cal F},4}^{(2)}$ are given by its OPE variables as (78), where $q^{2}$ can be specified as $1$, and $\{T,S\}$ need to be chosen as the kinematics belong to the $34\to 12(-q)$ process. For general $x=1/2$ and $10<\delta<$ 1, the inequalities will be
@@
In addition, $|T^{2}|$ should be far less than $|\delta|$ to capture the limit behavior of $\delta$. We can specify them to be
$T^{2}=-10^{-16}\,,\qquad S^{2}=-\delta^{-1}-\frac{1}{2}\,.$ (90)
@@
Then, the form factor result in the triple-collinear limit is obtained by setting $\{T,S\}$ as

### v3

@@ 2.1 Stress-tensor supermultiplet and the form factor
$-\theta_{\alpha}^{+a}\epsilon^{\alpha\beta}\theta_{\beta}^{b}{\rm tr}\Big{(}\psi^{+\gamma}_{(a}\psi^{+}_{b)\gamma}-g_{\text{YM}}\sqrt{2}[\phi^{+C}_{(a},\bar{\phi}_{C\,+b)}]\phi^{++}\Big{)}$
$-{4\over 3}(\theta^{+})^{3\,a}_{\alpha}{\rm tr}\Big{(}F_{\beta}^{\alpha}\psi_{a}^{+\beta}+ig_{\text{YM}}[\phi_{a}^{+B},\bar{\phi}_{BC}]\psi^{C\alpha}\Big{)}+{1\over 3}(\theta^{+})^{4}\,{\cal L}(x)\ .$ (2)
@@ 2.2 Unitarity-cut method
We also use $\psi_{ABC}(p)\equiv\varepsilon_{ABCD}\psi^{D}(p)$ to denote the four-dimensional fermion field carrying the negative helicity. Hence, the supersymmetry is preserved in this scheme. [^1] We recall that in the four-dimensional cuts, the helicity sum can be performed by integrating over the Grassmann numbers of internal lines $\sum_{\textrm{physical states of}\ \ell}\xrightarrow{4-\text{dim}}\int\text{d}^{4}\eta_{\ell}$.
@@ 3.2 $D$-dimensional unitarity cuts
The $q^{2}$-cut channel shown in Figure 1(a) with external states as $(1^{+},2^{+},3^{\phi_{12}},4^{\phi_{12}})$ is given by [^2] We use the abbreviation $i^{+}$ to denote $i^{g_{+}}$, which carries the positive helicity.
@@
The $q^{2}$-cut channel involves only the cut scalar states. A more complicated case is the $s_{12}$-cut channel shown in Figure 1(b), which contains gluonic and fermionic contributions and provides additional checks for our result. We consider the external states as $(1^{+},2^{\phi_{12}},3^{\phi_{12}},4^{+})$, then the cut result is
@@
where the subscript $\Phi_{i}$ denotes the physical state associated with $p_{i}$. The possible physical states $(\Phi_{5},\Phi_{6})$ are $(g^{-},\phi_{34})$, $(\phi_{34},g^{-})$, $(\psi_{134},\psi_{234})$, and $(\psi_{234},\psi_{134})$. The most complicated terms are $(\psi_{134},\psi_{234})$ and $(\psi_{234},\psi_{134})$ that involve the fermionic tree-level results. In this case, the two tree-level building blocks can be also computed by Feynman diagrams to capture the $D$-dimensional information, as
@@
where $p_{56}$ denotes $p_{5}+p_{6}$, and the $D$-dimensional gamma traces ${\bf tr}$ defined in (18) keep the $D$-dimensional information of the loop momenta, as we discussed in Section 2. We mention here that it is practically convenient to compare them with the cut integrand numerically.333 We give a brief comment on the numerical method for $D$-dimensional kinematics. Any massless momentum can be expressed as $\ell=(\ell^{0},\ell^{0}\hat{n}_{D-1})$, where $\ell^{0}$ is the zeroth component, and $\hat{n}_{D-1}=(\cos\theta_{1}\,,\ \ldots\,,\ \sin\theta_{1}\cdots\sin\theta_{D-2})$ is a $(D-1)$-dimensional unit vector. We consider the $D>4$ to capture the information beyond the four-dimension. This form takes the advantage that the momentum conservation equations and on-shell conditions are linear for $\ell^{0}$ (except the three-point case). Then it is convenient to apply the numerical evaluation with specific dimensions such as $D=6$, which is enough for the two-loop four-point case.
@@ 4.1 Summary of the two-loop integrand
$n_{2}=$ $s_{123}\left[s_{12}\text{tr}_{-}(1,\ell_{7},4,3)-s_{12\ell_{1}}\text{tr}_{-}(1,\ell_{7},4,3)+\text{tr}_{-}(1,\ell_{7},4,3,\ell_{1},2)\right]$
@@ 4.2 More details in the unitarity cuts
This unitarity cut has only contribution from $N_{1}$, since only the topology in Figure 2(a) contributes to this cut.777 The integrand based on four-dimensional cuts has been considered in Gopalka:2023doh . There seems no contribution of topology associated with $N_{1}$ in the result of Gopalka:2023doh , which would be inconsistent with the cut of Figure 3 even in the four-dimension. The $N_{1}$ part has the non-trivial contribution in the IBP reduced form and to the remainder function as will be considered in the next section.
@@
The $D$-dimensional version of the unitarity cut in Figure 3 with external states as $(1^{+},2^{+},3^{\phi_{12}},4^{\phi_{12}})$ is
@@
The $D$-dimensional unitarity cut as shown as Figure 3 with external states as $(1^{+},2^{+},3^{\phi_{12}},4^{\phi_{12}})$ involves the two-gluons cut states as
${\cal F}^{(2)}_{4}\big{|}_{\text{cut-(b)}}=$$\sum_{\{\Phi_{i}\}}{\cal A}_{4}^{(0)}(1^{+},2^{+},5^{\Phi_{5}},6^{\Phi_{6}}){\cal A}_{4}^{(0)}(3^{\phi_{12}},4^{\phi_{12}},\bar{7}^{\Phi_{\bar{7}}},\bar{8}^{\Phi_{\bar{8}}}){\cal F}_{2}^{(0)}(\bar{6}^{\Phi_{\bar{6}}},\bar{5}^{\Phi_{\bar{5}}},8^{\Phi_{8}},7^{\Phi_{7}})$
$=$ ${\cal A}_{4}^{(0)}(1^{+},2^{+},5^{-},6^{-}){\cal A}_{4}^{(0)}(3^{\phi_{12}},4^{\phi_{12}},\bar{7}^{\phi_{34}},\bar{8}^{\phi_{34}}){\cal F}_{2}^{(0)}(\bar{6}^{+},\bar{5}^{+},8^{\phi_{12}},7^{\phi_{12}})\,.$ (41)
Note that the external states of the cut legs for $p_{5}$ and $p_{6}$ must be gluon states $g^{-}$, and then the cut legs for $p_{7}$ and $p_{8}$ can only be scalar states $\phi_{12}$.
Another two-double-cut example similar to the one-loop case is shown in Figure 3 with external states as $(1^{\phi_{12}},2^{\phi_{12}},3^{+},4^{+})$
@@
Finally, the triple cut will provide the most stringent crosscheck for the integrand result. We find that all triple cuts pass directly given the result derived by the two-double cuts. Consider for example the cut shown in Figure 3 with external states as $(1^{+},2^{+},3^{\phi_{12}},4^{\phi_{12}})$:
@@ 5.1 The IBP reduction of the form factor
$\mathbb{I}_{i}=\sum_{a}c_{i,a}\mathbb{I}_{i,a}^{\text{IBP-input}}\,,$ (45)
where $\mathbb{I}_{i,a}^{\text{IBP-input}}$ are the IBP-input integrals, defined as
$\mathbb{I}_{i,a}^{\text{IBP-input}}=\prod_{k}D_{i,k}^{-\alpha_{i,a,k}}\,,$ (46)
where $D_{i,k}$ is the $k$-th propagator of $i$-th topology (including irreducible numerators).
@@
The IBP-input integrals are ready to be reduced to a finite set of master integrals by using IBP relations. We perform the IBP reduction by the program Kira Maierhofer:2017gsa ; Klappert:2020nbg in combination with Fermat Lewis and FireFly Klappert:2019emp ; Klappert:2020aqs . The coefficients of master integrals in the result can be simplified significantly for the half-BPS form factor if choosing the master integrals to have a uniform transcendental degree. The related UT master integrals, which will be denoted by $I^{(\ell),\text{UT}}$, are given in Abreu:2020jxa ; Canko:2020ylt ; Abreu:2021smk ; Abreu:2023rco .
@@
The parity-odd part is defined by the terms that are proportional to $\varepsilon(1234)$. Note that the $\varepsilon(1234)$ variable appears in both the factor $B$ and the definition of the parity-odd integral $I_{\text{Pen}}^{(1)}$. We mention that $I_{\text{Pen}}^{(1)}$ contributes only to the $\mathcal{O}(\epsilon)$ order, and so does ${\cal G}_{2}^{(1)}$ and ${\cal I}_{{\cal F},4}^{(1)}\big{|}_{\text{odd}}$.
@@ 5.2 Two-loop remainder function
where $i,j=1,2,3,4$ and note that only five of them are independent. We find that the remainder function in terms of the Chen iterated path integrals contains only 34 function letters $\{m_{i}\}$ which are the same as the symbol letters. We have checked that our result is consistent with the symbol expression of the same form factor obtained via symbol bootstrap in Dixon:2022xqh .
@@ 5.3 Evaluating the remainder function
Figure 8: Evaluating the remainder function for a physical scattering process. For the figures in the first row, the lower-left linear boundary and the upper-right curve boundary correspond separately to the outgoing angle $\theta=0$ and $\theta=\pi$. The two boundaries intersect symmetrically at $(0,1-m_{q}^{2})$ and $(1-m_{q}^{2},0)$ correspond separately to the soft limit with $p_{1}\to 0$ and $p_{2}\to 0$, where $m_{q}=1/10$.
@@
$s_{12}=P^{2}\,,\quad s_{13}=-\frac{1}{6}\,,\quad s_{14}=-\frac{1}{18}\,,\quad s_{23}=-\frac{1}{3}\,,\quad s_{24}=-\frac{1}{9}\,,\quad s_{34}=\frac{5}{3}\,,$ (70)
where $P^{2}$ will tend to 00 in the collinear limit. Then, we evaluate the form factor at this kinematics point by selecting a series of small $P^{2}$ and compare the numerical result with the known three-point form factor result. We observe a trend of convergence in the limit process when the parameter $P^{2}$ close to 00 but not equal to 00. The numerical result with $P^{2}=10^{-16}$ is
${\cal R}_{{\cal F},4}^{(2)}=4.991676051765813+0.05741704625875813i\,.$ (71)
The ${\cal R}_{{\cal F},3}$ is evaluated at the corresponding kinematics point at
$s_{12}^{\prime}=-1/2\,,\quad s_{23}^{\prime}=5/3\,,\quad s_{13}^{\prime}=-1/6\,,$ (72)
with $p_{1}^{\prime}=p_{1}+p_{2}$, $p_{2}^{\prime}=p_{3}$ and $p_{3}^{\prime}=p_{4}$. We find the difference as
${\cal R}_{{\cal F},4}^{(2)}-{\cal R}_{{\cal F},3}^{(2)}=-2.0\times 10^{-13}-3.1\times 10^{-14}i\,.$ (73)
Furthermore, in Figure 8, the remainder function at the lower-left linear boundary corresponds to $\theta=0$. We have checked that this is consistent with taking the collinear limit $p_{1}\parallel p_{2}$. For example, choosing the parameters $x_{1}=x_{2}=99/200+\delta$ and $\delta=10^{-16}$, the result is
${\cal R}_{{\cal F},4}^{(2)}=-5.398065883227598-3.949636854837923i\,.$ (74)
The corresponding kinematics of ${\cal R}_{{\cal F},3}^{(2)}$ are
@@
with $p_{1}^{\prime}=p_{1}+p_{2}$, $p_{2}^{\prime}=p_{3}$ and $p_{3}^{\prime}=p_{4}$, and the difference is
${\cal R}_{{\cal F},4}^{(2)}-{\cal R}_{{\cal F},3}^{(2)}=4.6\times 10^{-7}-6.6\times 10^{-10}i\,.$ (76)
The above two examples also show that the speed of convergence is not the same under different parameterizations.
@@
The lightlike form factor (with $q^{2}=0$) for the stress-tensor supermultiplet has been obtained using the master-integral bootstrap method in Guo:2022qgv , which is shown to have the directional dual conformal (DDC) symmetry and depends on only the three conformal ratios
@@
where $\delta\to 0$ and which satisfies $q^{2}=\delta$. Then the form factor is evaluated at this kinematics point with $\delta=10^{-16}$, as
${\cal R}_{{\cal F},4}^{(2)}=102.1176907432574-4.0\times 10^{-12}i\,,$ (79)
and the difference with respect to the lightlike form factor ${\cal R}_{{\cal F},4}^{(2),{\rm LL}}$ is
${\cal R}_{{\cal F},4}^{(2)}-{\cal R}_{{\cal F},4}^{(2),{\rm LL}}=-2.2\times 10^{-11}+4.0\times 10^{-12}i\,.$ (80)
In the next section, we will show that the DDC symmetry emerges also at the integrand level.
@@ 5.4 Triple-collinear limit
$u_{2}=\left[1+T^{2}+\frac{S^{2}[S_{2}T_{2}(1+f_{2}^{2})+f_{2}(1+S_{2}^{2}+T^{2}+T_{2}^{2})]}{f_{2}S_{2}^{2}}\right]^{-1}\,,$
@@
We note that the triple-collinear limit $T\to 0$ for parametrization (81) corresponds to $p_{4}\parallel p_{1}\parallel p_{2}$. Taking triple-collinear limit $p_{4}\parallel p_{1}\parallel p_{2}$ for the form factor in the kinematic region $34\to 12(-q)$ can be mapped to the triple-collinear limit $p_{1}\parallel p_{2}\parallel p_{3}$ for the form factor in the kinematic region $41\to 23(-q)$. The latter, following the above discussion, should be related to the triple-collinear limit ${\hat{p}}_{1}\parallel{\hat{p}}_{2}\parallel{\hat{p}}_{3}$ of six-point amplitude in the physical region ${\hat{4}}{\hat{1}}\to{\hat{2}}{\hat{3}}{\hat{5}}{\hat{6}}$. This correspondence is illustrated in Figure 9. In the following, we will compare the amplitude in the physical region ${\hat{4}}{\hat{1}}\to{\hat{2}}{\hat{3}}{\hat{5}}{\hat{6}}$ and the triple-collinear limit $p_{4}\parallel p_{1}\parallel p_{2}$ for the form factor in the kinematic region $34\to 12(-q)$ which should be equivalent up to a cyclic shift of momenta.
@@
It is notable that the full analytic remainder ${\cal R}_{{\cal A},6}^{(2)}$ in the physical region ${\hat{1}}{\hat{4}}\to{\hat{2}}{\hat{3}}{\hat{5}}{\hat{6}}$ is still not yet available in the literature. To apply our check, we note that the amplitude has been studied in a singular configuration, which corresponds to the limit in which a hexagonal Wilson loop develops a self-crossing Dixon:2016epj . In this self-crossing limit, the cross ratios are highly constrained, for example, for the process ${\hat{1}}{\hat{4}}\to{\hat{2}}{\hat{3}}{\hat{5}}{\hat{6}}$ one has
@@
Then the Mandelstam variables for ${\cal R}_{{\cal F},4}^{(2)}$ are given by its OPE variables as (81), where $q^{2}$ can be specified as $1$, and $\{T,S\}$ need to be chosen as the kinematics belong to the $34\to 12(-q)$ process. For $x=1/2$ and $10<\delta<$ 1, the inequalities will be
@@
In addition, to achieve the triple-collinear limit, $|T^{2}|$ should be far less than $|\delta|$. We can specify them to be
$T^{2}=-10^{-16}\,,\qquad S^{2}=-\delta^{-1}-\frac{1}{2}\,,$ (92)
where $-\frac{1}{2}$ in $S^{2}$ is a random choice, since the form factor in the triple-collinear limit should be independent of $S$.
@@
Then, the form factor result in the triple-collinear limit is obtained by solving for $\{T_{2},S_{2},f_{2}\}$ using (83), with $\{T,S\}$ specified as
@@
Note that the value of $S^{2}$ is a random choice in the allowed kinematic region, since the form factor in the triple-collinear limit should be independent of $S$.